Consumer spends all his or her income on food (X) and clothing (y). The current prices are p1 10 birr and P2 5birr and he or she maximizes utility by purchasing 20 units of X and 50 units of y.find how much he or she will purchase if she or he spends the entire income on x

1
Expert's answer
2022-04-18T15:17:31-0400

The consumer's income is:

I = 10*20 + 5*50 = 450 birr.

If she or he spends the entire income on x, then:

Qx = 450/10 = 45 units.